1. The demand and supply function for a good I start given as Q=3_2p and Q=20+2p.What will be the equilibrium price?
1
Expert's answer
2021-05-05T14:05:55-0400

"Qdemanded=Qsupplied"

"Qsupply=20-2p"

"Qdemanded=3-2p"

"3-2p=20+2p"

"-2p-2p=20-3"

"\u22124p\/-4=17\/\u22124"

"P=-4.25"

Equilibrium price= -4.25